An aluminum conductor (assume PEC) is located at 0 < ? < 1.6 m moves with velocity 2?⃗, m/s in a magnetic field given whose value is given by ? = 10 cos(???+ ) Wb/m2 where ? is a constant. Calculate the induced voltage in V.
Answer #1

In the above question the aluminium conductor is moving in a time varying magnetic field so we cannot directly apply the formula (E=B*L*V) .

So we need to apply faraday's law of electromagnetic induction [ E=d(phi)/dt ] and then solve the numerical.

Where (PHI=B*A) ,here B as well as A both are changing so we will use chain rule of differentiation while diffrentiating.